OpenStudy (hahahahaha):

Write x2 − 6x + 7 = 0 in the form (x − a)2 = b, where a and b are integers. (x − 4)2 = 3 (x − 3)2 = 2 (x − 2)2 = 1 (x − 1)2 = 4

OpenStudy (danjs):

You know the completing the square method?

OpenStudy (hahahahaha):

thats what im confused on

OpenStudy (danjs):

as is, it will not factor to that form, it can be (x-6)*(x-1)=0, but they want that (x-a)^2 = b ...

OpenStudy (danjs):

x^2 - 6x + 7 =0 move 7 over... \[x^2 - 6x = -7\]

OpenStudy (danjs):

To complete the square on that left side, you can add to both sides of the equation, half of that b term squared... in ax^2 + bx + c \[\large (\frac{ b }{ 2 })^2\]

OpenStudy (danjs):

i can show you where that comes from if you want, or you can just remember that..

OpenStudy (hahahahaha):

isnt that using completing the square and using the vertex or something?

OpenStudy (hahahahaha):

nvm i really dont know what im talking about XD

OpenStudy (danjs):

yeah if you take that half of b squared and add it on to the thing, it will become a square that factors to (x - a)^2 \[\large a*x^2+b*x\] \[\large ax^2 + bx + (\frac{ b }{ 2 })^2\]

OpenStudy (danjs):

you have a -6 there, so half of that squared is \[\large (\frac{ -6 }{ 2 })^2 = 9\] x^2 - 6x = -7 add that to both sides of the equation \[\large x^2 - 6x + 9 = -7 + 9\]

OpenStudy (hahahahaha):

a=1 b=-6 c=7 right? now i just plug it in?

OpenStudy (hahahahaha):

oh you already did it

OpenStudy (danjs):

As is it is not a perfect square, first move the 7 over to the other side to get x^2 - 6x = -7 now add on half of (-6) squared, or 9 to both sides x^2 - 6x + 9 = -7 + 9 \[\large x^2 - 6x + 9 = 2\] now it is a square that will factor \[\large (x - 3)^2 = 2\]

OpenStudy (hahahahaha):

ohhhh ok

OpenStudy (hahahahaha):

thanks for helping^

OpenStudy (danjs):

welcome, yeah just remember half b squared , to add on and complete the square..

OpenStudy (hahahahaha):

alright:)